OpenStudy (anonymous):

Use the second derivative test to find all relative extrema. ..... h(t) = t-4 times the root of (t+1).

OpenStudy (anonymous):

which one is h(t)? \(\large h(t)=t-4\sqrt{t+1} \) or \(\large h(t)=(t-4)\sqrt{t+1} \)

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

did you find h'(t) ?

OpenStudy (anonymous):

d'(t) = 1 - 2 / root of t+1

OpenStudy (anonymous):

write it as t - 4(t + 1)^0.5 h'(t) = 1 - 2(t + 1) ^(-0.5) equate this to 0 to find turning points on the graph

OpenStudy (anonymous):

i made a right mess of that - must be tired ill try again 0 = 1 - 2 / sqrt(t+1) sqrt(t+1) = 2 t+ 1 = 4 t = 3 = thats it

OpenStudy (anonymous):

now you need to know if thats a minimum or maximum second derivate h"(t) = 1 / ( t+ 1)^ 1.5 plug t = 3 into this gives a positive value therefore h(t) has a minimum value at t = 3 and that value is found by plugging t = 3 into t - 4 sort(t + 1) = 3 - 4 sort(4) = -5 so u have a minimum at (3,-5)